### v1

@@ Abstract
This paper develops small vision language models to understand visual art, which, given an art work, aims to identify its emotion category and explain this prediction with natural language. While small models are computationally efficient, their capacity is much limited compared with large models. To break this trade-off, this paper builds a small emotional vision language model (SEVLM) by emotion modeling and input-output feature alignment. On the one hand, based on valence-arousal-dominance (VAD) knowledge annotated by psychology experts, we introduce and fuse emotional features derived through VAD dictionary and a VAD head to align VAD vectors of predicted emotion explanation and the ground truth. This allows the vision language model to better understand and generate emotional texts, compared with using traditional text embeddings alone. On the other hand, we design a contrastive head to pull close embeddings of the image, its emotion class, and explanation, which aligns model outputs and inputs. On two public affective explanation datasets, we show that the proposed techniques consistently improve the visual art understanding performance of baseline SEVLMs. Importantly, the proposed model can be trained and evaluated on a single RTX 2080 Ti while exhibiting very strong performance: it not only outperforms the state-of-the-art small models but is also competitive compared with LLaVA 7B after fine-tuning and GPT4(V). Code is available at here.
@@ 1 Introduction
To make the output emotion explanations more emotional, we borrow the idea of valence-arousal-dominance (VAD) modeling: each word in the explanation text is represented by a 3-dim vector defined in the VAD dictionary [23]. Because the VAD scheme provides rich psychological descriptions of the text, we 1) fuse the VAD text features with the classic text features to improve the model input, and 2) design a VAD head to enforce the output text explanation to have similar VAD vectors with those of the ground truth explanation. The two measures allow the language model to better understand and output emotional texts.
@@ 4.1 Baseline Structure: A Small Language Model
Overview. As shown in Fig. 2 (a), our baseline has three inputs: An art image, an emotion classification prompt $M$: ‘the emotion is _’, and the ground truth test explanation text $X$. The test explanation is used as both input and output in training and output only in inference. Outputs of the baseline are the predicted emotion category and language explanation $X$ of the prediction. Our system is composed of an image encoder, a text encoder, and a language decoder.
@@ 4.2 Proposed Improvements
Here $\alpha$ is hyper-parameter, whose impact is evaluated in Sec. 5.3.
@@ 5.2 Main Evaluation
Table 2: Comparing SEVLM with LLaVA 7b after fine-tuning (denoted as LLaVA-FT). Our model is very competitive in most metrics, especially ACC and EA which are important indicators, except that LLaVA-FT has significantly higher diversity (lower Unique score). Moreover, our model is 42 times smaller than LLaVA-FT.
Dataset
Method
Backbone
@@
[…]
v1.0
$*{}^{*}$ [18]
CLIP-Vicuna
61.8 60.0 55.5 30.7 16.4 9.0 14.3 30.7 84.3
SEVLM (ours)
CLIP-GPT2
65.6 62.1 54.2 30.3 16.4 9.2 13.9 30.4 62.0
v2.0
$*{}^{*}$ [18]
CLIP-Vicuna
38.4
39.3 53.2 30.6 16.9 9.5 14.3 30.7 84.5
SEVLM (ours)
CLIP-GPT2
44.2 42.6 51.8 30.4 17.2 10.1 13.9 30.4 63.6
@@
[…]
Figure 6: Ablation study of emotion encoder on ArtEmis v1.0 test set. ‘w/o Emotion Encoder’ denotes that the VAD vectors extracted from the emotion dictionary are directly set as emotion features (in Eq. 2 ) without encoding them by the emotion encoder.
Figure 6: Ablation study of emotion encoder on ArtEmis v1.0 test set. ‘w/o Emotion Encoder’ denotes that the VAD vectors extracted from the emotion dictionary are directly set as emotion features (in Eq. 2 ) without encoding them by the emotion encoder.
@@
Computational efficiency comparisons. Tab. 3 summarizes the comparisons with the baseline and LLaVA-FT. Compared with the baseline, our model adds marginally more parameters and longer inference speed. Both can be trained and tested using a single 2080Ti. Compared with LLaVA-FT, our model is 42 times smaller and consumes much less computational resources and less floating point operations (FLOPs), while having a higher inference speed. Together with the emotion understanding performance comparisons above, we have achieved a much better accuracy-efficiency trade-off for small vision language models.
Figure 8: Impact of hyperparameter $\alpha$ defined in Eq. 8 on the ArtEmis v2.0 combined validation set. Left to right: metrics ACC, EA, B4, M, and R are used. We set $\alpha=2$ for both ArtEmis v2.0 combined and ArtEmis v1.0 validation set sets.
@@ 5.3 Further Analysis
Detailed ablations and variants of the VAD emotion feature and contrastive head. We present results in Fig. 7 and Fig. 7, for the two components, respectively. For the VAD emotion feature, if we remove the emotion encoder (see Sec. 4.2.1 for its description), all metrics on the ArtEmis V1.0 test set become worse, as shown in Fig. 7. Moreover, the contrastive head use two types of negative samples as shown in Fig. 4; if we remove either of two, we also observe performance drop in Fig. 7. These ablation and variant studies further verify the reasonableness of our system.
Impact of hyperparameter $\mathcal{\alpha}$ on ArtEmis v2.0 dataset. As shown in Fig. 8, firstly, the performance of emotion metric ACC/EA and semantic metric M exhibit continuous improvement with increasing weight $\alpha$, reaching peak at $\alpha=2$. Secondly, the performance of semantic metric R/B4 improves with increasing weight $\alpha$, and it reaches the peak value in the interval $[0.1,2]$. With further increments in $\alpha$, these metrics experience performance drop. We choose $\alpha=2$ as a result. The same value is directly applied to ArtEmis v1.0 dataset.
@@ 5.4 Qualitative Results
Visualization In Fig. 1, we give visualization examples of different models on two datasets. We found that existing methods have insufficient performance in emotional understanding and can be divided into two types. In one case, when the emotion classification is correct, semantic errors will occur in the interpretation. For example, in Fig. 1 (a), SAT and NLX-GPT2 describe the wrong action ‘dancing’ and wrong object‘ horse’ respectively, while our method generates the correct explanation ‘the man is fighting with the man.’ The other is when the emotion prediction is wrong, the interpretation will not only have semantic errors but also a mismatch between the interpretation and the emotion category. For example, in Fig. 1 (c), the category ‘Contentment’ predicted by NLX-GPT2 is inconsistent with explanation, $i$.$e$., “the woman looks like she is about to cry.” These visualization comparison demonstrates the superiority of our approach for emotion understanding of artistic images.

### v2

@@ Abstract
This paper develops small vision language models to understand visual art, which, given an art work, aims to identify its emotion category and explain this prediction with natural language. While small models are computationally efficient, their capacity is much limited compared with large models. To break this trade-off, this paper builds a small emotional vision language model (SEVLM) by emotion modeling and input-output feature alignment. On the one hand, based on valence-arousal-dominance (VAD) knowledge annotated by psychology experts, we introduce and fuse emotional features derived through VAD dictionary and a VAD head to align VAD vectors of predicted emotion explanation and the ground truth. This allows the vision language model to better understand and generate emotional texts, compared with using traditional text embeddings alone. On the other hand, we design a contrastive head to pull close embeddings of the image, its emotion class, and explanation, which aligns model outputs and inputs. On two public affective explanation datasets, we show that the proposed techniques consistently improve the visual art understanding performance of baseline SEVLMs. Importantly, the proposed model can be trained and evaluated on a single RTX 2080 Ti while exhibiting very strong performance: it not only outperforms the state-of-the-art small models but is also competitive compared with LLaVA 7B after fine-tuning and GPT4(V). The code is available at https://github.com/BetterZH/SEVLM-code.
@@ 1 Introduction
To make the output emotion explanations more emotional, we borrow the idea of valence-arousal-dominance (VAD) modeling: each word in the explanation text is represented by a 3-dim vector defined in the VAD dictionary [23]. Because the VAD scheme provides rich psychological descriptions of the text, we 1) fuse the VAD text features with the classic text features to improve the model input [^1] In image captioning [2, 22, 34], language explanations are used both as input and output during training but only as output in inference., and 2) design a VAD head to enforce the output text explanation to have similar VAD vectors with those of the ground truth explanation. The two measures allow the language model to better understand and output emotional texts.
@@ 4.1 Baseline Structure: A Small Language Model
Overview. As shown in Fig. 2 (a), our baseline has three inputs: An art image, an emotion classification prompt $M$: ‘the emotion is _’, and the ground truth test explanation text $X$. Outputs of the baseline are the predicted emotion category and language explanation $X$ of the prediction. Our system is composed of an image encoder, a text encoder, and a language decoder.
@@ 4.2 Proposed Improvements
Here $\alpha$ is hyper-parameter, whose impact is evaluated in the supplementary material.
@@ 5.2 Main Evaluation
Table 2: Comparing SEVLM with LLaVA 7B after fine-tuning (denoted as LLaVA-FT). Our model is very competitive in most metrics, especially ACC and EA which are important indicators, except that LLaVA-FT has significantly higher diversity (lower Unique score). Moreover, our model is 37.5 times smaller than LLaVA-FT.
@@
v1.0 LLaVA-FT [18] CLIP-Vicuna 61.8 60.0 55.5 30.7 16.4 9.0 14.3 30.7 84.3
SEVLM (ours) CLIP-GPT2 65.6 62.1 54.2 30.3 16.4 9.2 13.9 30.4 62.0
v2.0 LLaVA-FT [18] CLIP-Vicuna 38.4 39.3 53.2 30.6 16.9 9.5 14.3 30.7 84.5
SEVLM (ours) CLIP-GPT2 44.2 42.6 51.8 30.4 17.2 10.1 13.9 30.4 63.6
@@
Computational efficiency comparisons. Tab. 3 summarizes the comparisons with the baseline and LLaVA-FT. Compared with the baseline, our model adds marginally more parameters and longer inference speed. Both can be trained and tested using a single 2080Ti. Compared with LLaVA-FT, our model is 37.5 times smaller and consumes much less computational resources and less floating point operations (FLOPs), while having a higher inference speed. Together with the emotion understanding performance comparisons above, we have achieved a much better accuracy-efficiency trade-off for small vision language models.
@@
Figure 6: Ablation study of emotion encoder on ArtEmis v1.0 test set. ‘w/o Emotion Encoder’ denotes that the VAD vectors extracted from the emotion dictionary are directly set as emotion features (in LABEL:eq:_emotion-enhanced_f_X ) without encoding them by the emotion encoder.
@@ 5.3 Further Analysis
Detailed ablations and variants of the VAD emotion feature and contrastive head. We present results in Fig. 7 and Fig. 7, for the two components, respectively. For the VAD emotion feature, if we remove the emotion encoder (see Sec. 4.2.1 for its description), all metrics on the ArtEmis V1.0 test set become worse, as shown in Fig. 7. Moreover, the contrastive head use two types of negative samples as shown in Fig. 4; if we remove either of two, we also observe performance drop in Fig. 7. These ablation and variant studies further verify the reasonableness of our system. More experimental results, e.g., the impact of hyperparameter $\alpha$ and more analysis with existing methods are included in the supplementary material.
@@ 5.4 Qualitative Results
Visualization. In Fig. 1, we give visualization examples of different models on two datasets. We found that existing methods have insufficient performance in emotional understanding and can be divided into two types. In one case, when the emotion classification is correct, semantic errors will occur in the interpretation. For example, in Fig. 1 (a), SAT and NLX-GPT2 describe the wrong action ‘dancing’ and wrong object ‘horse’ respectively, while our method generates the correct explanation ‘the man is fighting with the man.’ The other is when the emotion prediction is wrong, the interpretation will not only have semantic errors but also a mismatch between the interpretation and the emotion category. For example, in Fig. 1 (c), the category ‘Contentment’ predicted by NLX-GPT2 is inconsistent with explanation, i.e., ‘the woman looks like she is about to cry.’ These visualization comparison demonstrates the superiority of our approach for emotion understanding of artistic images. See more examples in our supplementary.
@@ 6 Conclusion
Acknowledgements
This work is supported by the National Natural Science Foundation of China (62272144, 72188101, 62020106007, and U20A20183), the Major Project of Anhui Province (202203a05020011), and the Fundamental Research Funds for the Central Universities (JZ2024HGTG0309).